OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played one-fourth as long as they did at the second party, which was one-half as long as the third party. If they played for 7 minutes at the first party, how many minutes did they play at the third party? three and one-half minutes fourteen minutes fifty-six minutes one hundred twelve minutes

OpenStudy (anonymous):

@Mertsj can you help

OpenStudy (mertsj):

Let x = minutes played at the third party. How long did they play at the second party?

OpenStudy (anonymous):

@Mertsj im confused

OpenStudy (mertsj):

It says at the second party they played half as long as at the third party.

OpenStudy (mertsj):

So if they played 10 minutes at the third party, how long did they play at the second party?

OpenStudy (anonymous):

@Mertsj um

OpenStudy (mertsj):

Now please don't tell me that you don't know what one half of 10 is.

OpenStudy (anonymous):

I think I know the answer to the problem its 14. omg I do lol

OpenStudy (anonymous):

@Mertsj or 3 1/2

OpenStudy (mertsj):

If they played x minutes at the third party, they played (1/2)x at the second party.

OpenStudy (anonymous):

@Mertsj so 3 1/2

OpenStudy (mertsj):

If they played (1/2)x minutes at the second party, they played (1/4)(1/2)x at the first party.

OpenStudy (mertsj):

(1/4)(1/2)x = (1/8)x

OpenStudy (mertsj):

So since we know they played 7 minutes at the first party, we an write: \[\frac{1}{8}x=7\]

OpenStudy (mertsj):

Solve for x

OpenStudy (anonymous):

is it 3 1/2 or not?

OpenStudy (anonymous):

Or 14

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

So you don't know how to solve the equation I posted?

OpenStudy (mertsj):

How do you solve this equation: \[\frac{1}{2}x=5\]